A 5.30 mL sample of an H3PO4 solution of unknown concentration is titrated with a 1.100×10-2 M NaOH solution. A volume of 7.32 mL of the NaOH solution was required to reach the equivalence point. What is the concentration of the unknown H3PO4 solution? Express your answer with the appropriate units.
Answer #1

The reaction between H3PO4 and NaOH is as follows

H3PO4 + 3 NaOH ------------------ Na3PO4 + 3 H2O

1 mole      3 mole

NaOH = 1.100 x10^-2M and 7.32mL

number of moles of NaOH = 1.100 x10^-2 x 0.00732L = 0.00008052 moles

number of moles of NaOH = 0.00008052 moles

according to equation

3 moles of NaOH = 1 mole of H3PO4

0.00008052 moles of NaOH = ?

                                         = 0.00008052 x 1 / 3 = 0.00002684 moles

number of moles of H3PO4 = 0.00002684 mole

volume of H3PO4 = 5.30 mL = 0.00530 L

Concentration of the solution = number of moles / volume in L

                                              = 0.00002684 / 0.00530 = 0.005064 mol/L

Concentration of H3PO4 = 0.005064 mol/L = 5.06 x10^-3M

Concentration of H3PO4 = 5.06 x10^-3M.